OpenStudy (sloppycanada):

The degree measure of one of two complementary angles is 30 less than twice that of the other. What are the degree measures of the angles? (Remember that complimentary angles add to be 90 degrees) 45 degrees and 45 degrees 30 degrees and 60 degrees 27 degrees and 63 degrees 40 degrees and 50 degrees

OpenStudy (welshfella):

if the angles are x any then we can write y = 2x - 30 also as the angles rae complementary x + y = 90 you can find the angles by solving this system of equations

OpenStudy (welshfella):

write the 2 equations in same form y = 2x - 30 y = 90 - x subtract:- 0 = 3x -120 3x = 120 what is x?

OpenStudy (sloppycanada):

x = 40

OpenStudy (welshfella):

right so what is y equal to?

OpenStudy (sloppycanada):

50.

OpenStudy (welshfella):

yea you got it The last option is the correct one.